Wednesday, August 31, 2005

Water Test 31/8/05

Tank pH Temp Cond
A1 8.4 21.4 1.17
A2 8.1 17.3 1.36
A4 7.9 17.4 1.07

T1 6.2 17.2 1.27
T2 6.1 17 1.27
T3 5.9 17.1 1.32
T4 5.9 16.9 1.27
T5 5.8 17.1 1.32

Results indicate...
* Change 50% of water to correct excessive pH in A1, A2, A4
** the need for the addition of Bi-Carb Soda to sump, re-circulation system T1-T5

Clean & Maintenance 31/8/05

All fish well. Cleaned ARC floor. Rinses Sump and risers. Backwashed. Will need major FULL clean by end of week. Fish feeding well.

Added 500gm Salt, 100gm Bi-Carb Soda to sump

New Reservoir for Hydroponics

Thanks to Mr Godwin who plumbed the 1000l tank to store backwashed water from our ARC recirculation system to be reused in hydroponics.

Monday, August 29, 2005


The Silver Perch quarentined died. It was frozen to examine later. No other fish appear to have problems.
1 * Dead jumping Trout (thanks to Daniel for the email report)

Saturday, August 27, 2005

Feed and Check 27/8/05

Fish fed well.
Silver Perch (adult) moved to Aquarium 3- isolated from others (quarantine) due to possible fin rot. Added salt to A3 provide high salt content salt bath to attempt cure.
Added 1kg of Salt to recirculation system through sump to combat any possibility of transfer (if fin rot is confirmed).

Friday, August 26, 2005

Check and Adjust Water Flow 26/8/05

Continued concern over Silver Perch (Adult) showing continued signs of 'fin rot'?. Sluggish, exposed flesh on Tail Fin (Caudal) & Anal Fin.
Adjusted water flow to individual tanks. Removed added ornaments to Aquarium 1?? Checked Murray Cod OK however looked stressed. Fed Australian Bass and Murray Cod (Blood Worm and Brine Shrimp). Will need to quarantine Silver Perch if condition worsens.
Image from
Swept floor to remove spilt food and debris to minimise rodents etc. Moved new aquarium to LAB to prepare for establishment of Marine Tank (brackish water) with Bream, Flathead, Mullet etc. as part of Yr9 MAT Mini- Projects.

John, Nick, Corey- Fed all.

Thursday, August 25, 2005

Water Test 25/8/05

Team (2) - Water Test

Team (1) - Measuring Weight/Mass of a Take-a-way plastic container of....
Trout Pellets- 378g
Dust & Crumble- 182g
Australian Native Pellets- 657g
Silver Perch Pellets- 110g
* Will need to re-measure when all containers have equal quantity.

** General Observations- Large Silver Perch showing odd behavior and damage to fins.

Wednesday, August 24, 2005

Clean 24/8/05

Rinsed Risers
Removed debris and excess food from bottom of Tanks 1-5
Backwashed Filter
Changed 30% of water in Aquarium 1
Topped up (50%) of Water in Aquarium 2 (unused at this time)
Tidied ARC and LAB
* great work by our general assistant in cutting grass, wipper snippering the ARC grounds and removing trusses.
** Rainwater Tank is DOWN to ONLY 1/3 (33%) of capacity
*** Construction Class is working on building a walkway to protect plumbing behind tanks. Thanks Guys
**** Mr Godwin has invented an automatic Fish Feeder and almost completed installation. This will ensure that our fish are fed JUST the right amount at JUST the right time. This will dramatically increase growth espicially in the Trout.

Corey & Jermaine- Fed all. Fed 'guppy' Gambusia, Mosquitofish to Australian Bass.

Tuesday, August 23, 2005

Water Test 23/8/05

Air Temperature ARC- 14.3
Temp. pH Conductivity
AQ 1 18.2 7.1 1.2 (Murray Cod)
AQ 4 10.9 6.6 1.07 (Porter's Creek Yabbies)

TANK 1-5 13.3 6.6 1.07 (Trout, Silver Perch, Australian Bass)

John, Nick, Corey, Saxon- No Blood Worms- Fed all. Wood pile gone, yard looks great, mowed today.

Friday, August 19, 2005

Clean 19/8/05

Cleaned Sump, did not flush and rinse. Backwashed system. Stored all food away from rodents to keep safe and dry. Created 'dust & crumble' using coffee grinder for Silver Perch fingerlings. Dropped off new waste paper baskets for oysters (biofilter) photos soon.

Need a major clean- still over feeding probably due to increased temperatures Trout eating less?
Tidied up ARC removed rubbish.


Another 500g of bicarb was added today. pH is still around 6.0

Wednesday, August 17, 2005

Chemical Adjustment 17/8/05

We only use 2 chemicals in the Aquaculture Recirculation System- Salt and Bi-Carb Soda.
500g of Bi-Carb Soda was added today. This, as indicated in previous postings is used to correct pH by increasing pH to our prefered operating levels of 6-7 pH. The Bi-Carb also has a buffering affect in minimising pH swing, hardness/calcium content of the water etc. The Bi-Carb and ALL chemicals (Salt) are added to the Sump so that it is gently/slowly combined with the recirculated water.

Tuesday, August 16, 2005

Silver Perch (not interested in 'guppies')

A week ago some 'guppies' were introduced into T4 (large Silver Perch) tank to entice feeding. So far the Silver Perch have ignored the 'guppies'? Are Silver Perch Omni, Herb or Carnivores?
Research and come back with comments.

Maintenance and Clean 16/8/05

Cleaned all 'risers'
Created vortex/whirlpool to move sediment to centre of individual Tanks and flushed.
Backwashed filter, cleaned primary filter basket, drained and rinsed sump.
Refilled Sump with Tank water, 250g of salt and restarted recirculation system
Change approx 1/3 of A1
Hosed floor
* Need to add 500g of Bicarb Soda

Monday, August 15, 2005

Future Focus

We intend to expand the ARC to cater for greater numbers of fish. We have applied for several grants and are about to submit applications for another 2. Details will be added to the Blog soon.

Spring is fast approaching and we MUST get our garden plots built and ready to sow. Plots will be allocated to Stages eg. Stage 1, 2, 3, Classes and subject areas.
The ARC will be establishing a Stawberry, Tomato and Lettuce hydroponic farm to reuse recirculated ARC water as well as a Pumpkin patch including 'Giant Pumpkins (Atlantic Giant Pumpkin).

Crops and animal will enable a revenue 'stream' and comlimentary crops/food for the ARC. It is hoped that we can establish an interschool pumpkin growing comp.

Worms, chooks etc. are on their way.

Water Test 15/8/05

Tank/Aquarium Temp. pH Conduct.

A1 17.8 8.0 0.92
A2 11.0 7.7 1.41
A4 11.2 7.5 0.82

T1 12.6 4.8 0.97
T2 12.6 4.7 1.0
T3 12.6 4.6 1.02
T4 12.5 4.4 1.02
T5 12.6 4.4 0.92

* Tank 1-5 are seriously low in pH, with the water too acid. We have purchased some plastic covered metal mesh waste baskets which will contain Oyster Shells. The baskets will be located in the BioFilters and as acidic water passes over the shells the carbonate will be dissolved which will raise the pH and buffer the water against rapid pH swings. In the mean time we will add slowly 500 g of Bicarb Soda and retest.
** Aquariums will need at least 1/3 of the water exchanged to balance and refresh the water.

Wednesday, August 10, 2005

Completed Netting of T1

Hopefully the new net will stop our jumping Trout. Mr Heise and the whole class came up and fed the fish. New A4 tank with Yabbies and guppies doing great. Large Silver Perch in T4 seem to prefer a much reduced flow of water from the recirculation system...

* Lost another 'jumping' Trout

Tuesday, August 09, 2005

Extending the 'farm'

A recirculation aquaculture system produces nutrient rich water as waste.
We are looking to create a companion hydroponic garden to use the nutrient rich water to grow strawberries, lettuce and tomatoes.
To substitute commercial pelletised feed and support the diet of native species we are developing a worm farm to provide a natural food source.
The school produces a significant amount of waste (garbage) each and every school day. This includes: - paper which we will investigate to shred and compost - food scraps which can also be composted AND used to feed chooks. A local egg producer is closing in Jan/Feb 2006 and we have approached them to supply us with birds and equipment. This can be used as a revenue source for the ARC
We aim to develop a herbal tea concoction that will combine nutrient rich recirculation water and 'secrete ingredients to produce a microbiotic liquid fertiliser to grow pumpkins, also as a revenue source AND maybe to grow some extraordinary sized pumpkins to compete with other schools and communities. WHAT DO YOU THINK?

Mulberry Tree

After a very short month of no leaves the tree has sprouted new growth ready for new fruit. Adding value to the ARC.

* Irrigation: Although somewhat drought-resistant, mulberries need to be watered in dry seasons. If the roots become too dry during drought, the fruit is likely to drop before it has fully ripened.
* Fertilization: Mulberries generally thrive with minimal fertilization. An annual application of a balanced fertilizer such as 10:10:10 NPK will maintain satisfactory growth.

Clean and Net Construction

Purchased some polypipe to form a hoop to suspend netting around Tank 1 to prevent Trout from jumping. Began construction.
Cleaned all 'risers'
Placed 'sock' filter over inlet pipes to sump.
Created vortex/whirlpool to move sediment to centre of individual Tanks and flushed.
Backwashed, cleaned primary filter, drained and rinsed sump.
Added Tank water and restarted recirculation system.
Fed fish with John and Corey. Corey brought in small Yabbies and created Yabbie Tank in A4. Also some guppies(to be identified- maybe Mosquito Fish). John and Corey will get more Yabbies and will try to catch some Eels.
Hosed floor
Emptied garbage bin to school bin
* Lost another 'jumping' Trout

Monday, August 08, 2005

Maintenance and Classwork 8/8/05

We NEED rain...
Water Test including town tap water to check levels in case we need to use town water. We pride ourselves in supporting ZERO Town Water usage however it has been almost 5 weeks since our last significant rain. The tanks is well under half. Approach. 4000l. If we were to use Town water we would need to leave the water 'sit' for a least 10 days to allow chemicals to dissipate.

Water Temperature
tank1 tank2 tank3 tank4 tank5
13.0 13.0 13.0 13.4 13.0
aquarium1 aquarium2
17.5 11.0 (turned heater off- no fish)

tank1 tank2 tank3 tank4 tank5
6.5 6.5 6.5 6.5 7.4

ms (salt content)
tank1 tank2 tank3 tank4 tank5
0.87 0.87 0.87 0.87 0.87
aquarium1 aquarium2
1.00 1.41

tap:0.24 ???? What is this reading- we would expect approx 7.0 for pH and 0.0 for ms

Cut more squares of shade cloth to expand the capacity of the biofilter
Cleaned 'first rinse' device
Tested automatic feeder- GREAT!

Continued development of Badge Maker PowerPoint Presentation
Term review/s- Bookwork, Network Folder, eMail- Unfinished Work N's

Feed 8/8/05

To ensure maximum growth and care students will feed and check the ARC Period/s 1 & 6. Thanks to John, Nicolas, Jermaine and Corey and for their help.

Friday, August 05, 2005

Fish Feed and Clean

Fish continue to feed well, even the Trout even though air temps have exceeded expected and average daily temps. Asisted by John, Nicolas and Corey. These students will take responsibility for morning feeding (P1). This will enable us to 'ramp up' the feeding of the Trout- just 3 months to harvest!
Cleaned and rinsed 'risers', swept and hosed ARC. Backwashed filter, cleaned primary filter basket and topped up sump. Used the sock filter to catch uneaten food and 'fish poo' when cleaning risers. Worked well, still need to ensure that socks will not come loose in the sump and block the filter!

Wednesday, August 03, 2005

Total Clean 3/8/05

Full Clean and Maintenance.
Flushed and rinsed all Tanks
Drained sump, flushed and refilled
Remove and cleaned all Risers
Backwashed and Rinsed Filter
Swept and hosed ARC floor
Morning Feed of fish

* with help from John and Joel